# Milky Way-est: Cosmological Zoom-in Simulations with Large Magellanic Cloud and Gaia–Sausage–Enceladus Analogs

### v1

@@ 4.2 Subhalo Mass Functions
The left panel of Figure 7 shows mean SHMFs for Milky Way-est (purple) and Symphony MW (black). SHMFs are measured using present-day virial mass and normalized to each host’s virial mass at $z\approx 0$ (or at $z=0$ for Symphony MW). Milky Way-est systems host a mean number of $77\pm 13$ subhalos with $M_{\mathrm{sub}}/M_{\mathrm{host}}>1.4\times 10^{-4}$ (or $M_{\mathrm{sub}}\gtrsim 10^{8}~{}{M_{\odot}}$), versus $60\pm 13$ for Symphony MW, where uncertainties represent the host-to-host standard deviation. Thus, at low masses, the amplitude of the mean Milky Way-est SHMF is $22\%$ higher than in Symphony MW, while the SHMF slopes are consistent. This difference is statistically significant given the $\approx 3\%$ Poisson error on our stacked SHMF measurement and represents a $\approx 1.5\sigma$ difference relative to the host-to-host scatter. At higher masses ($M_{\text{sub}}/M_{\text{host}}\gtrsim 10^{-2}$), LMC analogs enhance Milky Way-est SHMFs relative to Symphony MW. This follows because the highest-$V_{\mathrm{peak,sub}}$ subhalos in Symphony MW are roughly one order of magnitude less massive than our LMC analogs, on average (see Section 3.2 and Figure 3).
@@ 4.2.1 LMC-associated Subhalos
The SHMF enhancement in Milky Way-est is partially due to subhalos accreted with our LMC analogs. To demonstrate this, the dashed lines in the top-left panel of Figure 7 show the contribution to the total SHMFs from the LMC analogs in Milky Way-est, and from the highest-$V_{\mathrm{peak,sub}}$ subhalos in Symphony MW. We identify LMC-associated subhalos within the LMC analog’s virial radius at the timestep before LMC accretion into the MW virial radius, following Nadler et al. (2020). Comparing the mean SHMFs, we find that Milky Way-est LMC analogs contribute a mean of $9\pm 6$ subhalos down to our $M_{\mathrm{sub}}/M_{\mathrm{host}}>1.4\times 10^{-4}$ resolution limit, where uncertainties again represent the host-to-host standard deviation. Meanwhile, subhalos associated with the highest-$V_{\mathrm{peak,sub}}$ systems in Symphony MW only contribute $3\pm 3$ subhalos down to the same limit. The dot-dashed line in the bottom-left panel of Figure 7 shows the SHMF ratio excluding LMC (or highest-$V_{\mathrm{peak,sub}}$)-associated subhalos and confirms that the difference is not entirely due to subhalos within the LMC’s virial radius when it accretes.
@@
Thus, $\approx 50\%$ of the difference between Milky Way-est and Symphony MW SHMF amplitudes can be directly attributed to LMC-associated subhalos. To assess the robustness of this result, we repeat our calculation using LMC-associated subhalos within $2{R_{\rm vir}}$ of the LMC before that region overlaps with each Milky Way-est host’s virial radius. This increases the mean abundance of LMC-associated subhalos in Milky Way-est from $9$ to $18$ and explains the remaining difference in SHMF amplitudes.
@@
Figure 8: Fraction of subhalos with $M_{\mathrm{sub}}>1.2\times 10^{8}~{}{M_{\odot}}$ contributed by Milky Way-est LMC analogs (purple circles) and Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos (gray squares) versus $M_{\mathrm{peak,sub}}/M_{\mathrm{host}}$ for the LMC analog or highest-$V_{\mathrm{peak,sub}}$ subhalo. Filled (unfilled) gray squares show Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos that pass (do not pass) our LMC analog $V_{\mathrm{max,sub}}$ and $a_{\mathrm{infall}}$ cuts. The dashed black line shows the one-to-one relation.
@@
Figure 8 shows the fraction of each host’s total subhalo population with $M_{\mathrm{sub}}>1.2\times 10^{8}~{}{M_{\odot}}$ contributed by Milky Way-est LMC analogs (purple circles) and Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos (gray squares). This fraction scales roughly linearly with $M_{\mathrm{peak,sub}}/M_{\mathrm{host}}$, which is reasonable given that subhalo abundance scales linearly with host mass (e.g., Nadler et al. 2023b). Subhalos associated with Milky Way-est LMC analogs constitute a mean of $(10\pm 6)\%$ of the total subhalo population down to $M_{\mathrm{sub}}>1.2\times 10^{8}~{}{M_{\odot}}$, versus $(4\pm 4)\%$ for systems associated with Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos, where uncertainties represent the host-to-host standard deviation. This $\approx 1\sigma$ difference follows because, on average, our LMC analogs are more massive and accrete more recently than Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos. In particular, Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos within our LMC analogs’ range of $a_{\mathrm{infall}}$ and $V_{\mathrm{max,sub}}$ contribute $(7\pm 5)\%$ of the total subhalo population, on average (filled gray squares in Figure 8), which is still slightly lower than the LMC-associated subhalo contribution in Milky Way-est. It will therefore be interesting to study how SHMF enhancement depends on our LMC analogs’ orbits (D’Souza & Bell, 2021; Barry et al., 2023).
@@ 4.2.2 GSE-associated Subhalos
Subhalos accreted with our GSE analogs may also contribute to the difference in SHMFs (e.g., see Bose et al. 2020). By identifying the number of subhalos within the GSE analog’s virial radius at the time of accretion, we find a mean of $3\pm 3$ surviving GSE-associated subhalos with $M_{\mathrm{sub}}>1.2\times 10^{8}$ in Milky Way-est hosts at $z\approx 0$, on average, where uncertainties represent the host-to-host standard deviation. This contribution is small compared to the LMC-associated subhalos discussed above. However, we caution that it is difficult to robustly track subhalos associated with our GSE analogs throughout their mergers and over the subsequent $\approx 10~{}\mathrm{Gyr}$ using Rockstar and Consistent-Trees. We leave a more detailed analysis of GSE-associated subhalos using particle tracking, which can follow subhalos much longer (e.g., Mansfield et al. 2023), to future work.
Figure 9: Projected maps of subhalo counts stacked over all Milky Way-est (left) and Symphony MW (right) hosts. Distances are measured in units of each host’s virial radius before stacking. Subhalo counts are normalized by the total number of subhalos in each host, such that each pixel shows the fraction of subhalos in a given projected area. Subhalo populations are rotated in three dimensions such that the center of each Milky Way-est LMC analog (white circles) or Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalo (white squares) lie along the $x$-axis. White dashed rings correspond to $1/3~{}R_{\mathrm{vir,host}}$, $2/3~{}R_{\mathrm{vir,host}}$, and $R_{\mathrm{vir,host}}$.
@@ 4.2.3 Summary
We have shown that Milky Way-est systems host $22\%$ more subhalos with $M_{\mathrm{sub}}\gtrsim 10^{8}~{}{M_{\odot}}$, on average, than Symphony MW hosts that are not selected to satisfy our GSE and LMC analog criteria. GSE-associated subhalos contribute negligibly to this difference. Subhalos within the virial radius of the LMC at the time of LMC infall explain half of the difference, and subhalos within $2{R_{\rm vir}}$ of the LMC when that region is entirely outside of the host explain the remaining difference. We have confirmed that Milky Way-est and Symphony MW SHMFs are similar at $a=0.8$, consistent with these results. Finally, we note that secondary host properties are unlikely to affect this interpretation; in particular, although host concentration correlates with subhalo abundance in general (Mao et al., 2015), the difference in SHMF amplitude persists when excluding Symphony MW hosts outside of the Milky Way-est host concentration range.
@@ 4.3 Subhalo Radial Distributions
The enhancement of Milky Way-est inner radial distributions is partially driven by our LMC analogs. At $z\approx 0$, each LMC analog is roughly $50~{}\mathrm{kpc}$ from the host center; for a typical Milky Way-est host with $R_{\mathrm{vir,host}}=300~{}{\rm kpc}$, this translates to $r/R_{\mathrm{vir,host}}\approx 0.2$. Within this distance, LMC-associated subhalos constitute $27\%$ of the mean Milky Way-est subhalo abundance, compared to $10\%$ within the entire virial radius, on average. At larger distances, the contribution of LMC-associated subhalos decreases. This can be understood because (before the LMC system’s first pericenter) LMC-associated subhalos are mostly contained within the LMC analog’s virial radius of $\approx 150~{}{\rm kpc}$, which extends to $\approx 200~{}{\rm kpc}$ from the host center.
Thus, subhalos directly associated with the LMC contribute to the enhancement of the inner radial distribution in Milky Way-est relative to Symphony MW but do not entirely explain the difference between the suites. The dot–dashed line in the bottom-right panel of Figure 7 shows the radial distribution ratio excluding LMC (or highest-$V_{\mathrm{peak,sub}}$)-associated subhalos, and (as for the SHMF) again confirms that the difference is not entirely due to the LMC. Although GSE-associated subhalos are expected to preferentially be found in the inner regions of the host (e.g., Bose et al. 2020), our tests in Section 4.2 indicate that these systems do not significantly contribute to present-day Milky Way-est subhalo abundances given our analysis tools. It is therefore unlikely that the remaining radial distribution difference is explained GSE-associated subhalos. As for the SHMF, future work that studies the origins of subhalos that contribute to the inner radial distribution difference will be useful.
@@ 4.4 Spatial Anisotropy of Subhalo Populations
Milky Way-est LMC analogs accrete recently, accompanied by their own subhalos (see Section 4.2). Because we analyze most LMC analogs at or near their first pericenter, we expect that LMC-associated subhalos are still clustered near their LMC analog host at $z\approx 0$ (e.g., D’Souza & Bell 2021), inducing spatial anisotropy in the subhalo population in excess of that for an “average” MW-mass halo. To assess this, we measure the spatial anisotropy of subhalo populations by calculating the fraction of subhalos in the hemisphere that points from the host center to the LMC analog at the $z\approx 0$ snapshot in each Milky Way-est simulation; an anisotropy measure of $50\%$, therefore, corresponds to an isotropic subhalo distribution. In Symphony MW, we use the highest-$V_{\mathrm{peak,sub}}$ subhalo at $z=0$ to define the preferred direction.
We find that subhalos are preferentially located in the present-day direction of our LMC analogs and that this effect is strongest in the inner regions of Milky Way-est hosts. In particular, the mean anisotropy of Milky Way-est subhalo populations is $58\%$ within $1/3~{}R_{\mathrm{vir,host}}$, $56\%$ within $2/3~{}R_{\mathrm{vir,host}}$, and $54\%$ within $R_{\mathrm{vir,host}}$. In Symphony MW, we observe little to no anisotropy, with $51$ to $52\%$ of all subhalos found in the highest-$V_{\mathrm{peak,sub}}$ subhalo’s hemisphere at all radii. For both suites, the host-to-host standard deviation of these measurements is $\approx 5\%$. These differences between Milky Way-est and Symphony MW are statistically significant, given the $\approx 3\%$ Poisson error on our stacked anisotropy measurements, and represent a $\approx 1\sigma$ difference relative to the host-to-host scatter.
@@ 5.1.1 Comparison to Previous Work
Using two MW-like simulations with LMC analogs from the original Mao et al. (2015) Symphony MW suite, Nadler et al. (2020) demonstrated that spatial anisotropy of subhalo and satellite galaxy populations driven by the LMC system is necessary to fit the observed MW satellite population. The LMC analogs studied in Nadler et al. (2020) respectively have sub-to-host halo mass ratios of $0.1$ and $0.2$, so they enhance the SHMF, inner radial distribution, and spatial anisotropy by a typical amount relative to our LMC analog sample (see Figure 8). Our LMC-associated subhalo results are also consistent with Lu et al. (2016), who found a similar enhancement of the satellite stellar mass function and spatial anisotropy due to satellites accreted with LMC analogs in the original Symphony MW suite (Mao et al., 2015).
We note that Barry et al. (2023) report even larger LMC-related SHMF enhancements using hydrodynamic zoom-in simulations run with the Feedback in Realistic Environments (FIRE) code. The central galaxies in these simulations preferentially disrupt subhalos that accrete earlier (Garrison-Kimmel et al., 2017; Nadler et al., 2018), which may enhance the relative contribution to the SHMF from LMC-associated subhalos. Future work will explore this effect by resimulating Symphony MW and Milky Way-est hosts with embedded disk potentials (Y. Wang et al., in preparation).

### v2

@@ 4.2 Subhalo Mass Functions
The left panel of Figure 7 shows mean SHMFs for Milky Way-est (purple) and Symphony MW (black). SHMFs are measured using present-day virial mass and normalized to each host’s virial mass at $z\approx 0$ (or at $z=0$ for Symphony MW). Milky Way-est systems host a mean number of $77\pm 13$ subhalos with $M_{\mathrm{sub}}/M_{\mathrm{host}}>1.4\times 10^{-4}$ (or $M_{\mathrm{sub}}\gtrsim 10^{8}~{}{M_{\odot}}$), versus $60\pm 13$ for Symphony MW, where uncertainties represent the host-to-host standard deviation. Thus, at low masses the amplitude of the mean Milky Way-est SHMF is $22\%$ higher than in Symphony MW, while the SHMF slopes are consistent. This difference is statistically significant given the $\approx 3\%$ Poisson error on our stacked SHMF measurement and represents a $\approx 1.5\sigma$ difference relative to the host-to-host scatter. At higher masses ($M_{\text{sub}}/M_{\text{host}}\gtrsim 10^{-2}$), LMC analogs enhance Milky Way-est SHMFs relative to Symphony MW. This follows because the highest-$V_{\mathrm{peak,sub}}$ subhalos in Symphony MW are roughly 1 order of magnitude less massive than our LMC analogs, on average (see Section 3.2 and Figure 3).
@@ 4.2.1 LMC-associated Subhalos
The SHMF enhancement in Milky Way-est is partially due to subhalos accreted with our LMC analogs. To demonstrate this, the dashed lines in the top-left panel of Figure 7 show the contribution to the total SHMFs from the LMC analogs in Milky Way-est, and from the highest-$V_{\mathrm{peak,sub}}$ subhalos in Symphony MW. We identify LMC-associated subhalos within the LMC analog’s virial radius at the time step before LMC accretion into the MW virial radius, following Nadler et al. (2020). Comparing the mean SHMFs, we find that Milky Way-est LMC analogs contribute a mean of $9\pm 6$ subhalos down to our $M_{\mathrm{sub}}/M_{\mathrm{host}}>1.4\times 10^{-4}$ resolution limit, where uncertainties again represent the host-to-host standard deviation. Meanwhile, subhalos associated with the highest-$V_{\mathrm{peak,sub}}$ systems in Symphony MW only contribute $3\pm 3$ subhalos down to the same limit. The dotted–dashed line in the bottom-left panel of Figure 7 shows the SHMF ratio excluding LMC (or highest-$V_{\mathrm{peak,sub}}$)-associated subhalos and confirms that the difference is not entirely due to subhalos within the LMC’s virial radius when it accretes.
@@
Thus, $\approx 50\%$ of the difference between Milky Way-est and Symphony MW SHMF amplitudes can be directly attributed to LMC-associated subhalos. We also repeat our calculation using LMC-associated subhalos within $2{R_{\rm vir}}$ of the LMC before that region overlaps with each Milky Way-est host’s virial radius. This increases the mean abundance of LMC-associated subhalos in Milky Way-est from 9 to 18 and explains the remaining difference in SHMF amplitudes, indicating that the remaining excess subhalos in Milky Way-est also fell in from a spatial region associated with the LMC.
@@
Figure 8: Fraction of subhalos with $M_{\mathrm{sub}}>1.2\times 10^{8}~{}{M_{\odot}}$ contributed by Milky Way-est LMC analogs (purple circles) and Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos (gray squares) vs. $M_{\mathrm{peak,sub}}/M_{\mathrm{host}}$ for the LMC analog or highest-$V_{\mathrm{peak,sub}}$ subhalo. Filled (unfilled) gray squares show Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos that pass (do not pass) our LMC analog $V_{\mathrm{max,sub}}$ and $a_{\mathrm{infall}}$ cuts. The dashed black line shows the one-to-one relation.
@@
Figure 8 shows the fraction of each host’s total subhalo population with $M_{\mathrm{sub}}>1.2\times 10^{8}~{}{M_{\odot}}$ contributed by Milky Way-est LMC analogs (purple circles) and Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos (unfilled and filled gray squares). This fraction scales roughly linearly with $M_{\mathrm{peak,sub}}/M_{\mathrm{host}}$, which is reasonable given that subhalo abundance scales linearly with host mass (e.g., Nadler et al. 2023b). Subhalos associated with Milky Way-est LMC analogs constitute a mean of $(10\pm 6)\%$ of the total subhalo population down to $M_{\mathrm{sub}}>1.2\times 10^{8}~{}{M_{\odot}}$, versus $(4\pm 4)\%$ for systems associated with Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos, where uncertainties represent the host-to-host standard deviation. This $\approx 1\sigma$ difference follows because, on average, our LMC analogs are more massive and accrete more recently than Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos. In particular, Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalos within our LMC analogs’ range of $a_{\mathrm{infall}}$ and $V_{\mathrm{max,sub}}$ contribute $(7\pm 5)\%$ of the total subhalo population, on average (filled gray squares in Figure 8), which is still slightly lower than the LMC-associated subhalo contribution in Milky Way-est. It will therefore be interesting to study how SHMF enhancement depends on our LMC analogs’ orbits (Patel et al., 2020; D’Souza & Bell, 2021; Barry et al., 2023).
@@ 4.2.2 GSE-associated Subhalos
Subhalos accreted with our GSE analogs may also contribute to the difference in SHMFs (see, e.g., Bose et al. 2020). By identifying the number of subhalos within each GSE analog’s virial radius at the time of accretion, we find a mean of $3\pm 3$ surviving GSE-associated subhalos with $M_{\mathrm{sub}}>1.2\times 10^{8}$ in Milky Way-est hosts at $z\approx 0$, on average, where uncertainties represent the host-to-host standard deviation. This contribution is small compared to the LMC-associated subhalos discussed above. However, we caution that it is difficult to robustly track subhalos associated with our GSE analogs throughout their mergers and over the subsequent $\approx 10~{}\mathrm{Gyr}$ using Rockstar and Consistent-Trees. We leave a more detailed analysis of GSE-associated subhalos using particle tracking, which can follow subhalos much longer (e.g., Mansfield et al. 2023), to future work.
Figure 9: Projected maps of subhalo counts stacked over all Milky Way-est (left panel) and Symphony MW (right panel) hosts. Distances are measured in units of each host’s virial radius before stacking. Subhalo counts are normalized by the total number of subhalos in each host, such that each pixel shows the fraction of subhalos in a given projected area. Subhalo populations are rotated in three dimensions such that the center of each Milky Way-est LMC analog (white circles) or Symphony MW highest-$V_{\mathrm{peak,sub}}$ subhalo (white squares) lie along the $x$-axis. White dashed rings correspond to $1/3~{}R_{\mathrm{vir,host}}$, $2/3~{}R_{\mathrm{vir,host}}$, and $R_{\mathrm{vir,host}}$.
@@ 4.2.3 Summary
We have shown that Milky Way-est systems host $22\%$ more subhalos with $M_{\mathrm{sub}}\gtrsim 10^{8}~{}{M_{\odot}}$, on average, than Symphony MW hosts that are not selected to satisfy our GSE and LMC analog criteria. GSE-associated subhalos contribute negligibly to this difference. Subhalos within the virial radius of the LMC at the time of LMC infall explain half of the difference, and subhalos within $2{R_{\rm vir}}$ of the LMC when that region is entirely outside of the host explain the remaining difference. We have confirmed that Milky Way-est and Symphony MW SHMFs are similar at $a=0.8$ ($z=0.25$), consistent with these results. Finally, we note that secondary host properties are unlikely to affect this interpretation; in particular, although host concentration correlates with subhalo abundance in general (Mao et al., 2015), the difference in SHMF amplitude persists when excluding Symphony MW hosts outside of the Milky Way-est host concentration range.
@@ 4.3 Subhalo Radial Distributions
The enhancement of Milky Way-est inner radial distributions is partially driven by our LMC analogs. At $z\approx 0$, each LMC analog is roughly $50~{}\mathrm{kpc}$ from the host center; for a typical Milky Way-est host with $R_{\mathrm{vir,host}}=300~{}{\rm kpc}$, this translates to $r/R_{\mathrm{vir,host}}\approx 0.2$. Within this distance, LMC-associated subhalos constitute $27\%$ of the mean Milky Way-est subhalo abundance, compared to $10\%$ within the entire virial radius, on average. At larger distances, this contribution of LMC-associated subhalos decreases. This can be understood because (before the LMC system’s first pericenter) LMC-associated subhalos are mostly contained within the LMC analog’s virial radius of $\approx 150~{}{\rm kpc}$, which extends to $\approx 200~{}{\rm kpc}$ from the host center.
Thus, subhalos directly associated with the LMC contribute to the enhancement of the inner radial distribution in Milky Way-est relative to Symphony MW but do not entirely explain the difference between the suites. The dotted–dashed line in the bottom-right panel of Figure 7 shows the radial distribution ratio excluding LMC (or highest-$V_{\mathrm{peak,sub}}$)-associated subhalos, and (as for the SHMF) again confirms that the difference is not entirely due to the LMC. We note that subhalos within $2{R_{\rm vir}}$ of each LMC analog when that region is outside of the host account for most of the remaining difference within $100~{}{\rm kpc}$. Additionally, although GSE-associated subhalos are expected to preferentially be found in the inner regions of the host (e.g., Bose et al. 2020), our tests in Section 4.2 indicate that these systems do not significantly contribute to present-day Milky Way-est subhalo abundances given our analysis tools. It is therefore unlikely that the remaining radial distribution difference is explained by GSE-associated subhalos. As for the SHMF, future work that studies the origins of subhalos which contribute to the inner radial distribution difference will be useful.
@@ 4.4 Spatial Anisotropy of Subhalo Populations
Milky Way-est LMC analogs accrete recently, accompanied by their own subhalos (see Section 4.2).
Because we analyze most LMC analogs at or near their first pericenter, we expect that LMC-associated subhalos are still clustered near their LMC analog host at $z\approx 0$ (e.g., D’Souza & Bell 2021). This effect would induce spatial anisotropy in the subhalo population in excess of that for an “average” MW-mass halo, consistent with LMC-induced clustering in angular momentum space reported by previous studies (e.g., Garavito-Camargo et al. 2021b, 2023). To assess this, we measure the spatial anisotropy of subhalo populations by calculating the fraction of subhalos in the hemisphere that points from the host center to the LMC analog at the $z\approx 0$ snapshot in each Milky Way-est simulation; an anisotropy measure of $50\%$, therefore, corresponds to an isotropic subhalo distribution. In Symphony MW, we use the highest-$V_{\mathrm{peak,sub}}$ subhalo at $z=0$ to define the preferred direction.
We find that subhalos are preferentially located in the present-day direction of our LMC analogs and that this effect is strongest in the inner regions of Milky Way-est hosts. In particular, the mean anisotropy of Milky Way-est subhalo populations is $58\%$ within $1/3~{}R_{\mathrm{vir,host}}$, $56\%$ within $2/3~{}R_{\mathrm{vir,host}}$, and $54\%$ within $R_{\mathrm{vir,host}}$. In Symphony MW, we observe little to no anisotropy, with $51\%$$52\%$ of all subhalos found in the highest-$V_{\mathrm{peak,sub}}$ subhalo’s hemisphere at all radii. For both suites, the host-to-host standard deviation of these measurements is $\approx 5\%$. These differences between Milky Way-est and Symphony MW are statistically significant, given the $\approx 3\%$ Poisson error on our stacked anisotropy measurements, and represent a $\approx 1\sigma$ difference relative to the host-to-host scatter.
@@ 5.1.1 Comparison to Previous Work
Using two MW-like simulations with LMC analogs from the original Mao et al. (2015) Symphony MW suite, Nadler et al. (2020) demonstrated that spatial anisotropy of subhalo and satellite galaxy populations driven by the LMC system is necessary to fit the observed MW satellite population. The LMC analogs studied in Nadler et al. (2020) have sub-to-host halo mass ratios of $0.1$ and $0.2$, respectively, so they enhance the SHMF, inner radial distribution, and spatial anisotropy by a typical amount relative to our LMC analog sample (see Figure 8). Our LMC-associated subhalo results are also consistent with Lu et al. (2016), who found a similar enhancement of the satellite stellar mass function and spatial anisotropy due to satellites accreted with LMC analogs in the original Symphony MW suite (Mao et al., 2015).
We note that Barry et al. (2023) reports even larger LMC-related SHMF enhancements using hydrodynamic zoom-in simulations run with the Feedback in Realistic Environments (FIRE) code. The central galaxies in these simulations preferentially disrupt subhalos that accrete earlier (Garrison-Kimmel et al., 2017; Nadler et al., 2018), which may enhance the relative contribution to the SHMF from LMC-associated subhalos. Future work will explore this effect by resimulating Symphony MW and Milky Way-est hosts with embedded disk potentials (Y. Wang et al. 2024, in preparation).